# Heterogeneous network estimation for single-cell transcriptomic data via a joint regularized deep neural network

### v1

@@ 2.2 Heterogeneous Analysis
In practice, the subgroup assignments $C_{i}$’s are not always observed. To accommodate the unknown cell heterogeneity, we propose adopting the K-means clustering with the Mahalanobis distance, which is defined as $dis_{ik}=\sqrt{\left(\boldsymbol{x_{i}}-\boldsymbol{o}_{k}\right)^{\mathrm{T}}\boldsymbol{\Sigma}_{k}^{-1}\left(\boldsymbol{x_{i}}-\boldsymbol{o}_{k}\right)}$, where $\boldsymbol{o}_{k}$ is the centroid of the $k$th subgroup and $\boldsymbol{\Sigma}_{k}$ is the covariance matrix for the $p$ genes. Instead of estimating $\boldsymbol{\Sigma}_{k}$, we consider estimating the precision matrix $\boldsymbol{\Theta}_{k}=\boldsymbol{\Sigma}_{k}^{-1}$ (which describes the conditional dependence between any two genes given the rest) directly with the network constructed using JRDNN. Specifically, $\hat{\boldsymbol{\Theta}}_{k}=(\hat{\Theta}_{k,jl})_{p\times p}$ is estimated with $\hat{\Theta}_{k,jl}=\hat{\Theta}_{k,lj}=1$ if $(j,l)\in\hat{E}_{k}$ and $\hat{\Theta}_{k,jl}=\hat{\Theta}_{k,lj}=0$ otherwise, for $j\neq l$, and $\hat{\Theta}_{k,ll}=1$, where $\hat{E}_{k}=\left\{(j,l):\hat{w}_{k,jl}\neq 0\textrm{ or }\hat{w}_{k,lj}\neq 0\right\}$ with $\hat{w}_{k,jl}$ being estimated using JRDNN. The proposed K-means clustering strategy effectively incorporates the network dependence among genes and takes into account heterogeneity attributed to both differential genes and gene relationships, leading to more accurate cell subgroup identification.
@@ 2.3 Computation
To obtain the final network estimation and cell subgroup identification results, we develop an iteration strategy (see Supplementary Algorithm S1) that conducts JRDNN and K-means clustering iteratively until convergence, where the subgroup memberships $C_{i}$’s obtained by the K-means clustering are delivered to JRDNN for joint estimation of multiple networks for the $K$ cell subgroups and the estimated networks $\hat{E}_{k}$ are used for computing the Mahalanobis distance in the K-means clustering. Specifically, for realizing the JRDNN step, we adopt the stochastic gradient descent (SGD) algorithm (see Supplementary Algorithm S2), which is implemented through the deep learning framework Pytorch in Python. To accommodate the discontinuity of the indicator function in (2.1), we adopt $1-e^{-\frac{w_{k,jl}^{2}}{\xi}}$ to approximate $1(w_{k,jl}\neq 0)$, where $\xi$ is a small positive constant and controls the goodness and smoothness of approximation. For realizing the K-means clustering step, we develop an iteration algorithm (see Supplementary Algorithm S3) that iteratively conducts a search for the cluster centroids and assigns the cluster memberships based on the Mahalanobis distance.
There are multiple tuning parameters in the proposed algorithm. For determining the number of subgroups $K$, we employ the silhouette coefficient (Shahapure & Nicholas, 2020), which maximizes the mean intra-cluster distance and minimizes the mean nearest-cluster distance simultaneously. For the three parameters in (2.1), instead of conducting the popular cross-validation, which is often time-consuming, we choose to consider some fixed values with the dependence on the numbers of subgroups and genes. Specifically, we set $\lambda_{1}=\frac{1}{p}$, $\lambda_{2}=\frac{1}{K\sqrt{p}}$, and $\lambda_{3}=\frac{1}{K\cdot\sqrt{d^{(1)}}}$, where $d^{(1)}=d_{0}^{(1)}+\sum_{k=1}^{K}d_{k}^{(1)}$ is the width of the first layer of JRDNN. For the hyperparameters of DNN, we choose a learning rate of 0.001, batch size of 50, activation function $tanh$, 200 epochs, and three hidden layers.
3 Simulation
3.1 Simulation Settings
In simulation, we consider 3,000 ($n=3,000$) cells from three subgroups ($K=3$), with measurements for 100 genes ($p=100$). For the subgroup sizes, two settings are considered, where the balanced design has 1,000 cells in each subgroup and the imbalanced design has 1,500, 900, and 600 cells in the three subgroups, respectively. For each subgroup, we consider the block-wise networks with ten unconnected sub-networks, which mimic the module-based properties of the practical biological networks. For each sub-network, two types of network structure are examined. The first is a star-chain structure, where the sub-network has one of the following patterns: (1) star pattern, there is one central node connected with all other nodes, and the non-central nodes are not connected with each other; (2) chain pattern, only the adjacent nodes are connected. The second is a scale-free structure, where the degree distribution follows a power law distribution with the maximum degree being four. The toy examples of these two types of network structures are provided in Supplementary Figure S1. The ratio of the common edges shared by three subgroups is $30\%$ for both types of network.
Based on the network, we consider the following mechanisms for generating the measurements $x_{ij}$’s. First, set a “core gene” $j_{1}$ in each sub-network, which is the central gene for the star pattern, the first gene for the chain pattern (denote the adjacent genes as $j_{1},j_{2},\cdots,j_{10}$), and the hub gene with the largest degree for the scale-free pattern. Second, we simulate $x_{ij_{1}}$ of the “core gene” $j_{1}$ from $N(0,1)$, $N(5,1)$, and $N(10,1)$ for the three subgroups, respectively. Then, for the star pattern, we simulate $x_{ij_{l}}=f\left(x_{ij_{1}}\right)+\epsilon_{i}$ for the other nine non-central genes, and for the chain pattern, we simulate $x_{ij_{l+1}}=f\left(x_{ij_{l}}\right)+\epsilon_{i}$ for $l=1,\cdots,9$ sequentially, where $f(\cdot)$ is a nonlinear function and $\epsilon_{i}\sim N(0,1)$ is a random noise. Three settings with from simple to complex nonlinear relationships for $f(\cdot)$ are investigated, including $f(x)=\sqrt[3]{x}$ (simple), $f(x)=0.25\exp(x)+\sqrt[3]{x}$ (medium), and $f(x)=3\sin(x)+0.25\exp(x)+\sqrt[3]{x}$ (complex). For the scale-free pattern, we first sort the genes $j_{1},\cdots,j_{10}$ by degree from largest to smallest and denoted $N_{j_{l}}$ as the node sets that have edges with $j_{l}$, then simulate $x_{ij_{l}}=\sum_{j_{k}\in N_{j_{l}}\text{ and }k<l}f_{k}(x_{ij_{k}})$ for $l=2,\cdots,10$ sequentially. Three settings for $f_{k}(\cdot)$’s are also considered, where all settings include polynomial term $x^{3}$ and interaction term $0.5x_{1}x_{2}$, and two additional terms $\sqrt{x}$ and $(x_{1}+x_{2})^{2}$ (simple), $0.2\exp(x)$ and $\log(x^{2})$ (medium), and $0.2\exp(x)$ and $\sin(x)$ (complex). We further simulate three dropout levels, where $[10\%,30\%]$ (low), $[20\%,40\%]$ (medium), and $[30\%,50\%]$ (high) elements of $x_{ij}$’s are replaced by zero. As a result, there are a total of 36 scenarios, covering a wide spectrum with various types of network structures, various levels of nonlinear relationships and dropouts, and both balanced and imbalanced sample distributions.
3.2 Competing Methods
In addition to the proposed JRDNN-KM, we consider ten competing methods, including BLGGM, JGNsc, JSEM, SpQN, Normalisr, GENIE3, locCSN, CS-CORE, SC3, and Seurat. Among them, BLGGM (Wu & Luo, 2022) is a Bayesian heterogeneous network analysis approach based on the zero-inflated Gaussian distribution, which can achieve multiple network estimation and subgroup identification simultaneously. Both JGNsc and JSEM conduct a heterogeneous network analysis with prior known subgroup information, where the common structures of networks among different cell subgroups are accommodated. Specifically, JGNsc (Dong et al., 2023) is a two-stage approach that first develops a Bayesian hybrid imputation method for accommodating zero-inflated data and then conducts joint graphical lasso for multiple network estimation. JSEM (Ma & Michailidis, 2016) conducts a joint analysis of multiple networks using node-wise linear regressions incorporating a group lasso penalty. SpQN, Normalisr, GENIE3, LocCSN, and CS-CORE were all originally developed for a single network analysis. Specifically, SpQN (Wang et al., 2022) is a two-stage approach that first conducts normalizing local distributions in correlation matrices to remove the mean-correlation relationship and then adopts the graphical lasso to refine network estimation. Normalisr (Wang, 2021) develops a normalization and statistical association testing framework to detect linear relationships with effective accommodation of the nonlinear confounders arising from library size. GENIE3 (Huynh-Thu et al., 2010) explores the conditional dependence based on the node-wise nonlinear regression and adopts an ensemble of tree-based machine learning algorithms. LocCSN (Wang et al., 2021) proposes a nonparametric investigation of the joint distribution of gene expression and can detect nonlinear correlations. CS-CORE (Su et al., 2023) is an unconditional correlation-based distribution-free approach capable of addressing the high sequencing depth variations and measurement errors in scRNA-seq data. For these approaches based on the prior known subgroup information, we consider the true cell subgroup memberships and conduct JGNsc and JSEM for joint network analysis and conduct SpQN, Normalisr, GENIE3, LocCSN, and CS-CORE separately for each subgroup. SC3 (Kiselev et al., 2017) and Seurat (Butler et al., 2018) are two popular heterogeneous analysis methods for single-cell transcriptomic data, without accounting for network estimation, where SC3 is a consensus clustering method and Seurat is a graph-based clustering based on the projected space from PCA. BLGGM, JGNsc, JSEM, and CS-CORE are realized using R codes from https://github.com/WgitU/BLGGM, https://github.com/meichendong/JGNsc, https://github.com/drjingma/JSEM, and https://github.com/ChangSuBiostats/CS-CORE, respectively. Normalisr is realized using Python codes from https://github.com/lingfeiwang/normalisr. In addition, SpQN, GENIE3, SC3, and Seurat are realized using R packages SpQN, GENIE3, SC3, Seurat, respectively, and locCSN is realized using Python package locCSN.
3.3 Examination of the number of subgroups
Given the candidate values $\{1,2,3,4,5\}$ of $K$, we take the simulation scenarios under the balanced design and low dropout rate as examples and examine the performance of the silhouette coefficient for selecting $K$, where two types of network structure and three levels of nonlinear relationships are considered. Specifically, the frequency at which a particular value is selected over 100 replicates is reported in Supplementary Table S1. In general, the silhouette coefficient consistently shows accurate identification of the true number of subgroups ($K=3$). As expected, the more complex network structure (scale-free) and nonlinear relationships make it slightly more difficult to identify the true value of $K$.
3.4 Evaluation of heterogeneity analysis and network estimation
In the downstream analysis, we set $K=3$ for all the methods to facilitate a more direct comparison and simulated 100 replicates under each scenario. The adjusted rand index (ARI) is adopted to assess the heterogeneity analysis performance, with values ranging from 0 to 1 and a larger value indicating a higher subgroup identification accuracy. In addition, we also employ Recall, Precision, and F1 score to evaluate the performance of network estimation, where the true and estimated edges are examined. Summarized ARI values and F1 scores over 100 replicates are provided in Figures 2 and 3, and the rest of the results are presented in Supplementary Figures S2 and S3. Here, ARI values for JGNsc, JSEM, SpQN, Normalisr, GENIE3, locCSN, and CS-CORE are not available since they are developed for network estimation and cannot identify unknown cell subgroups. In our study, we utilize true subgroup memberships for these methods. In addition, Recall, Precision, and F1 score are not available for SC3 and Seurat, as they do not conduct network estimation.
In Figure 2, JRDNN-KM demonstrate superior subgroup identification performance across all scenarios, especially as the level of sparsity intensifies and the complexity of nonlinear relationships increases. The benefits of JRDNN-KM that can effectively accommodate the nonlinear relationships among genes become increasingly apparent under these challenging conditions. Under scenarios with imbalanced samples, the subgroup identification accuracy of all methods diminishes, potentially due to the reduced sample size within some subgroups. However, JRDNN-KM still achieves higher ARI values. In general, BLGMM performs the second best, as it accommodates both the network structure and high sparsity of data. With the focus on the differential expressions only, SC3 and Seurat have the lowest ARI values.
In addition to the satisfactory subgroup identification performance, in Figure 3, JRDNN-KM consistently outperforms the alternative methods in network estimation accuracy, with larger values of F1 across all scenarios. Notably, as shown in Supplementary Figure S2, when the degree of nonlinearity increases, the Recall values of all methods become smaller, particularly for those reliant on linear assumptions, such as BLGGM, JGNsc, SpQN, and Normalisr. The proposed JRDNN-KM, which is based on deep neural networks, displays marked advantages as the nonlinear complexity escalates. Additionally, with an increase in zero inflation, all methods experience a decline in the F1 value. However, JRDNN-KM can maintain its superiority even amid increased dropout occurrences. It also shows robustness in situations with imbalanced sample sizes, with no significant performance differences observed compared to balanced settings. With the ability to accommodate nonlinear relationships, GENIE3 and locCSN maintain better performance than methods based on linearity assumptions. CS-CORE, being a distribution-free method that does not rely on distributional assumptions and may have the potential to handle nonlinearities, achieves good performance in nonlinear situations as well. The superiority of JRDNN-KM over these three methods directly supports the proposed DNN strategy for more effectively addressing complex nonlinear relationships and simultaneously taking into account the commonality among different subgroups. Although with the true subgroup memberships, the GMM-based JGNsc, JSEM, and SpQN lose effectiveness in the face of the complicated network structure, where JGNsc generally outperforms JSEM and SpQN by incorporating dropout. Overall, JRDNN-KM can achieve superior performance in both heterogeneity and network analysis, suggesting its practical usefulness in single-cell transcriptomic data analysis.
Figure 2: Summarized simulation results over 100 replicates with different methods. (A) and (B) Boxplots of ARI values under a balanced and imbalanced sample distribution, respectively. In each subfigure, two network structures (star-chain and scale-free), three levels of zero-inflation rate (low, medium, and high), and three levels of nonlinear relationships (simple, medium, and complex) are considered.
Figure 3: Summarized simulation results over 100 replicates with different methods. (A) and (B) of F1 scores under a balanced and imbalanced sample distribution, respectively. In each subfigure, two network structures (star-chain and scale-free), three levels of zero-inflation rate (low, medium, and high), and three levels of nonlinear relationships (simple, medium, and complex) are considered.
3.5 Sensitivity Analysis
We continue to investigate scenarios where only linear relationships exist among genes, leading to potential model misspecification for JRDNN-KM. Specifically, we consider the star-chain network structure with $f(x)=\beta x$, where $\beta\sim N(3,1)$. Under the balanced design, various numbers of cells ($n=3,000,6,000$, and $9,000$) and dropout levels are examined. The corresponding summary results are reported in Supplementary Figure S4. Under these relatively simpler scenarios, the four heterogeneity analysis methods all have satisfactory subgroup identification accuracy, especially for JRDNN-KM and BLGGM, which achieve average ARI values of 1. In terms of network estimation accuracy, with the limited sample size ($n=3000$), performance of JRDNN-KM is slightly weaker compared to strictly linear methods, such as BLGGM and JGNsc, due to the nonlinearity introduced by activation functions in DNN that may disrupt the inherent linearity. However, JRDNN-KM still outperforms the other nonlinear methods, GENIE3, locCSN, and CS-CORE, and the linear JSEM , SpQN, and Normalisr methods. When the sample size increases, JRDNN-KM demonstrates substantial improvements and even surpassed BLGGM and JGNsc, suggesting its effectiveness under the model misspecification.
Additionally, in Supplementary Table S2, we perform simulations with various levels of common information across different subgroups. Specifically, the balanced design with star-chain network, medium dropout rate, and medium nonlinear relationships are considered. It can be seen that for JRDNN-KM and JGNsc, which pay particular attention to the similarity among different subgroups, the increasing common information potentially facilitates a more accurate network estimation with a large value of F1. To further get deeper insights into the characteristics of JRDNN-KM, we also conduct ablation experiments. Specifically, under the scenarios with the balanced design, star-chain network, low dropout rate, and three levels of nonlinear relationships, we consider the proposed framework without the similarity regularization term or (and) the homogeneous hidden layers. Summary results are provided in Supplementary Table S3. It is observed that the proposed two techniques significantly contribute to both the heterogeneity analysis and network estimation. The significance is even more prominent under the scenarios with more complex nonlinear relationships.
@@ 4 Application
We analyze five publicly available single-cell transcriptomic datasets, on human lung adenocarcinoma (LUAD) cell lines (Tian et al., 2019), human peripheral blood mononuclear cells (PBMC) (Zheng et al., 2017), mouse embryonic stem cells (mESCs) (Kolodziejczyk et al., 2015), and mouse liver and uterus cells from the Mouse Cell Atlas (MCA) (Han et al., 2018). In each of these five datasets, every cell has been carefully labeled with cell or lineage markers. These labels have been served as gold standards, enabling objective evaluation of the cell subgroup identification performance. To control the quality of datasets, we follow the published studies and remove cells expressing less than 500 genes and genes expressed in less than 100 cells. Summary information of these five datasets is provided in Table 1, including the numbers of cells and genes, proportion of zeros, and subgroup information. These five datasets vary in sample size from small (mESCs) to large (PBMC) and are in data quality from high with a low zero-inflation rate (mESCs) to poor with zero-inflation rates exceeding 90% (PBMC). The sample distributions across different subgroups also show diversity, where the mouse uterus data has a highly imbalanced sample distribution.
All data are then subjected to batch effect correction using MNN (Haghverdi et al., 2018), normalized for library size using the total count normalization (TC) method (Cole et al., 2019), and filtered for 100 highly variable genes using the variance stabilizing transformation (vst) method in R package “Seruat” (Hafemeister & Satija, 2019).
Table 1: Summary information of the five real datasets
Dataset No. of cells No. of genes Proportion of zeros Subgroup information
LUAD 1,401 12,997 41.16% H2228 (461), H1975 (519),
HCC827 (421)
PBMC 12,418 8,560 91.12% CD19+B (2,341),
CD4+/CD25 Reg T (4,683),
CD56+NK (5,394)
mESCs 704 15,618 30.77% 2i (295), a2i (159), lif (250)
Mouse Liver 2,116 6,673 90.13% Dentritic (336),
Endothelial (621),
Erythroblast (199),
Kupffer (768), T (192)
Mouse Uterus 2,313 6,237 86.39% Endothelial (124),
Glandular (232),
Macrophage (141),
Muscle (223),
NK (135), Stromal (1,458)
@@ 4.1 JRDNN-KM leads to biologically sensible cell subgroups
The ARI values with JRDNN-KM as well as three competing methods are provided in Figure 4(A). The proportions of identified cell subgroups with JRDNN-KM in the true cell subgroups are reported in Figure 4(B) and the results with alternatives are reported in Supplementary Figure S5. It is observed that JRDNN-KM has superior or competitive performance across multiple datasets with satisfactory ARI values. Specifically, in the LUAD dataset, JRDNN-KM achieves an ARI value of 1.00. This performance markedly outstrips other alternative methods. Additionally, JRDNN-KM maintains higher performance metrics across diverse cellular contexts, including PBMC, mESCs, and mouse uterus cells, where the mouse uterus dataset has a highly imbalanced sample distribution with some subgroups having a small number of cells. For the mouse liver cell data, which has a high dropout rate (90.13%) and small sample size, JRDNN-KM performs slightly worse than BLGMM, probably due to the fact that the DNN strategy usually demands a relatively large sample size, but still behaves much better than SC3 and Seurat. Although there is a certain degree of mis-identification for some datasets, Figure 4(B) shows that each true subgroup is dominated by one of the subgroups identified with JRDNN-KM, further validating its utility in precision oncology and regenerative medicine. Similar to simulation, the two methods, SC3 and Seurat, that do not accommodate the network structure among genes, always have inferior cell subgroup identification accuracy.
@@
Figure 4: Heterogeneity analysis results on five real single-cell transcriptomic datasets. (A) ARI values with different methods. (B) Proportions of identified cell subgroups with JRDNN-KM in the true cell subgroups.

### v2

@@ 3.2 Heterogeneous Analysis
In practice, the subgroup assignments $C_{i}$’s are not always observed. To address the unknown cell heterogeneity in (Q1), we propose using K-means clustering with Mahalanobis distance, which is defined as $dis_{ik}=\sqrt{\left(\boldsymbol{x_{i}}-\boldsymbol{o}_{k}\right)^{\top}\boldsymbol{\Sigma}_{k}^{-1}\left(\boldsymbol{x_{i}}-\boldsymbol{o}_{k}\right)}$, where $\boldsymbol{o}_{k}$ is the centroid of the $k$th subgroup and $\boldsymbol{\Sigma}_{k}$ is the covariance matrix for the $p$ genes. Instead of estimating $\boldsymbol{\Sigma}_{k}$, we consider estimating the precision matrix $\boldsymbol{\Theta}_{k}=\boldsymbol{\Sigma}_{k}^{-1}$ (which describes the conditional dependence between any two genes given the rest) directly with the network constructed using JRDNN. Specifically, $\hat{\boldsymbol{\Theta}}_{k}=(\hat{\Theta}_{k,jl})_{p\times p}$ is estimated with $\hat{\Theta}_{k,jl}=\hat{\Theta}_{k,lj}=1$ if $(j,l)\in\hat{E}_{k}$ and $\hat{\Theta}_{k,jl}=\hat{\Theta}_{k,lj}=0$ otherwise, for $j\neq l$, and $\hat{\Theta}_{k,ll}=1$, where $\hat{E}_{k}=\left\{(j,l):\hat{w}_{k,jl}\neq 0\textrm{ or }\hat{w}_{k,lj}\neq 0\right\}$ with $\hat{w}_{k,jl}$ being estimated using JRDNN.
Here, binary $\Theta_{k,jl}$ acts solely as a dependency indicator, not a magnitude measure. This design aligns with Mahalanobis distance-based K-means clustering: K-means partitions data based on relative dissimilarity, while the Mahalanobis distance incorporating the precision matrix’s sparsity pattern discriminates local dependence from independence. This adaptation captures direct conditional relationships, ensuring clusters reflect similarity in the underlying graph-defined dependency network. The strategy accounts for heterogeneity in both gene expression levels and gene-gene relationships, thereby enabling more accurate cell subgroup identification.
@@ 3.3 Computation
To obtain the final network estimation and cell subgroup identification results, we implement an iterative strategy (Supplementary Algorithm S1) that alternates between JRDNN-based joint network estimation and Mahalanobis distance-based K-means clustering until convergence. Each iteration proceeds as follows: First, using current subgroup memberships $C_{i}$ derived from K-means clustering, we perform JRDNN estimation. Consistent with established practice for focusing on variable relationships rather than mean differences (Ma and Michailidis, 2016), we standardize the data within each subgroup during this step. The updated network estimate $\hat{\Theta}_{k}$ is then fed back into the K-means algorithm, which clusters the unstandardized data using this current estimate to compute Mahalanobis distances. For the JRDNN optimization, we minimize the penalized loss function (3.1) using stochastic gradient descent (SGD; Supplementary Algorithm S2) implemented via Python’s PyTorch framework. To address the discontinuity of the indicator function in (3.1), we approximate $1(w_{k,jl}\neq 0)$ with the smooth function $1-e^{-\frac{w_{k,jl}^{2}}{\xi}}$, where $\xi$ is a small positive constant and controls the approximation fidelity. The K-means clustering step employs our specialized algorithm (Supplementary Algorithm S3) that iteratively updates cluster centroids and reassigns memberships based on Mahalanobis distance. The complete parameter settings for all steps, examination of the memory and time usage of the proposed algorithm, and discussions on the convergence of the two-step iterative optimization are provided in Supplementary Section S2.
@@ 4 Application
We apply the proposed JRDNN-KM method to the five single-cell transcriptomic datasets in Section 2 and address questions (Q1)-(Q2). First, to control the quality of datasets, we follow the published studies and remove cells expressing less than 500 genes and genes expressed in less than 100 cells. All data are then subjected to batch effect correction using MNN (Haghverdi et al., 2018), normalized for library size using the total count normalization (TC) method (Cole et al., 2019), and filtered for 100 highly variable (HV) genes using the variance stabilizing transformation (vst) method in R package “Seruat” (Hafemeister and Satija, 2019).
Before further analysis, we perform a series of model checks for each cell type based on the gold-standard labels. This includes validating the JRDNN model’s capture of non-linear gene relationships through residual correlation analysis, and verifying the normality of residuals and the absence of outliers. As detailed in Section S3 of the Supplementary Materials, these checks confirm that the JRDNN method is appropriate for all five datasets.
In addition to the proposed JRDNN-KM, we consider ten competing methods, including BLGGM (Wu and Luo, 2022), JGNsc (Dong et al., 2023), JSEM (Ma and Michailidis, 2016), SpQN (Wang et al., 2022), Normalisr (Wang, 2021), GENIE3 (Huynh-Thu et al., 2010), locCSN (Wang et al., 2021), CS-CORE (Su et al., 2023), SC3 (Kiselev et al., 2017), and Seurat (Butler et al., 2018). Details for these competing methods and their implementation are provided in Section S4 of the Supplementary Materials.
@@ 4.1 JRDNN-KM leads to biologically sensible cell subgroups
In Figure 2(A), the Adjusted Rand Index (ARI) and Normalized Mutual Information (NMI) are used to evaluate subgroup identification performance. Both metrics range from 0 to 1, with higher values indicating greater accuracy. Note that ARI and NMI values are not available for JGNsc, JSEM, SpQN, Normalisr, GENIE3, locCSN, and CS-CORE, as these methods are designed for network inference and do not identify unknown cell subgroups. The proportions of cell subgroups identified by JRDNN-KM that correspond to the true subgroups are shown in Figure 2(B), with results for alternative methods provided in Supplementary Figure S4.
@@
Overall, JRDNN-KM demonstrates superior or competitive performance across multiple datasets, achieving satisfactory ARI and NMI values. This indicates that the proposed strategy effectively addresses (Q1), yielding biologically meaningful cell subgroups characterized by heterogeneous gene networks. Specifically, in the LUAD dataset, JRDNN-KM achieves both ARI and NMI values of 1.00. This performance markedly outstrips other alternative methods. Additionally, JRDNN-KM maintains higher performance metrics across diverse cellular contexts, including PBMC, mESCs, and mouse uterus cells, where the mouse uterus dataset has a highly imbalanced sample distribution with some subgroups having a small number of cells. For the mouse liver cell data, which has a high dropout rate (90.13%) and small sample size, JRDNN-KM performs slightly worse than BLGMM, but still behaves much better than SC3 and Seurat. Although there is a certain degree of mis-identification for some datasets, Figure 2(B) shows that each true subgroup is dominated by one of the subgroups identified with JRDNN-KM, further validating its utility in precision oncology and regenerative medicine. The two methods, SC3 and Seurat, that do not accommodate the network structure among genes, always have inferior cell subgroup identification accuracy.
@@
Figure 2: Heterogeneity analysis results on five real single-cell transcriptomic datasets. (A) ARI and NMI values with different methods. (B) Proportions of identified cell subgroups with JRDNN-KM in the true cell subgroups.
4.2 JRDNN-KM demonstrates high-accuracy alignment with well-established biological networks
To validate the accuracy of network reconstruction, we employ established biological network databases following the published studies (Chevalley et al., 2025; Pratapa et al., 2020). Specifically, we utilize three major databases: CORUM for protein complexes, STRING for protein-protein interactions (PPIs), and PerturbAtlas for genetic and chemical perturbations. Within STRING, we distinguish between two interaction types: PPI(F), which encompasses functional interactions, and PPI(P), which contains experimentally verified physical interactions. Given that the cell-type-specific information in these databases remains relatively limited, similar to many existing studies, we construct non-cell-type-specific networks here, which remain consistent across different cell types. For each of these four reference databases, we extract subnetworks by matching our 100 HV genes, with the resulting edge counts documented in Supplementary Table S3. Details regarding network construction and database specifications are available in Supplementary Section S5.
Figure 3: F1 scores of different methods evaluated against four reference networks across five datasets. Shapes represent methods, colors indicate reference networks, and horizontal lines show mean values across the four networks.
We evaluate network estimation performance using Recall, Precision, and F1 score by comparing the estimated edges against those from four established reference networks. F1 scores are summarized in Figure 3, while detailed Recall and Precision values are provided in Supplementary Figures S5 and S6. For the seven competing methods that require prior subgroup information, we supply the true subgroup memberships. Note that SC3 and Seurat are excluded from this comparison as they do not perform network estimation. The results show that our method achieves the highest F1 scores across all five datasets when compared to the CORUM-based and PerturbAtlas-based networks. Owing to the larger number of edges included in the PerturbAtlas-based network, which can lead to reduced recall, all methods exhibit relatively lower F1 scores in this context. For the two PPI networks, our method also delivers satisfactory performance. It attains the highest F1 scores in both PPI networks across the LUAD, mouse liver, and mouse uterus datasets. In the PBMC dataset with a high zero-inflation rate and the mESCs dataset with a limited sample size, our results on either PPI (F) or PPI (P) are slightly lower than those of BLGGM but still exceed the performance of other methods. In terms of average performance across all four networks, our method ranks first in four out of the five datasets (LUAD, PBMC, mouse liver, and mouse uterus) and second in the mESCs dataset, with performance nearly matching that of BLGGM. These findings validate the capability of the proposed network reconstruction framework in effectively addressing (Q2).